# Proper maps of annuli

### v1

@@ 5. Other remarks on proper maps of annuli
Let us put together some further minor remarks. Firstly, there are two other domains in $\mathbb{C}^{n}$ whose automorphism group is $U(n)$. There is the punctured ball $\mathbb{B}_{n}\setminus\{0\}$ and the complement of the closed unit ball $\mathbb{C}^{n}\setminus\overline{\mathbb{B}_{n}}$. For maps from a punctured ball to a punctured ball, the maps must extend through the origin by the Riemann extension theorem if $n=1$ or via the Hartogs phenomenon if $n>1$. We are then simply dealing with proper maps of balls that take the origin to the origin. Since the automorphism group of the ball is transitive, every spherical equivalence class of proper maps of balls contains a map taking 0 to 0. The case of the complement of the closed unit ball was studied in [helal-2025-proper]. In that case, for $n\geq 2$, the proper maps are polynomial sphere maps whose norm goes to infinity at infinity. In either case, the classification is different in these degenerate cases, and in both cases there are more maps than for annuli.
@@
The previous discussion suggests that the existence of a rational proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ of degree $d$ may imply a relationship between $r$, $R$, and $d$. For homogeneous maps as above we have $r^{d}\leq R$, with $H_{d}$ itself achieving the equality in this bound. Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general. If one also requires that $N$ is the embedding dimension ($N=N_{f}$), then more can be said. In this case, if $d=1$, then $R=r$. For higher degree, one could use juxtaposition to get a range. For example, if we juxtapose $H_{1}$ and $H_{d}$, then via the calculation in the introduction, we have that for $t\in(0,1)$, the map $\sqrt{1-t}\,H_{1}\oplus\sqrt{t}\,H_{d}$ has $R=\sqrt{(1-t)r^{2}+tr^{2d}}$, and so any $R$ in the range between $r^{d}$ and $r$ is achieved by these maps. So the question could be modified for the case when $N=N_{f}$ and degree $d>1$, to ask if $R$ equal to or bigger than $r$ is possible.

### v2

@@ 7. Other remarks on proper maps of annuli
Let us put together some further minor remarks. Firstly, there are two other domains in $\mathbb{C}^{n}$ whose automorphism group is $U(n)$. There is the punctured ball $\mathbb{B}_{n}\setminus\{0\}$ and the complement of the closed unit ball $\mathbb{C}^{n}\setminus\overline{\mathbb{B}_{n}}$. For maps from a punctured ball to a punctured ball, the maps must extend through the origin by the Riemann extension theorem if $n=1$ or via the Hartogs phenomenon if $n>1$. We are then simply dealing with proper maps of balls that take the origin to the origin. Since the automorphism group of the ball is transitive, every spherical equivalence class of proper maps of balls contains a map taking 0 to 0. The case of the complement of the closed unit ball was studied in [19]. In that case, for $n\geq 2$, the proper maps are polynomial sphere maps whose norm goes to infinity at infinity. In either case, the classification is different in these degenerate cases, and in both cases there are more maps than for annuli.
@@
The previous discussion suggests that the existence of a rational proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ of degree $d$ may imply a relationship between $r$, $R$, and $d$. For homogeneous maps as above we have $r^{d}\leq R$, with $H_{d}$ itself achieving the equality in this bound. Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems to be if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general. In Section˜6, we show that all degree 2 maps satisfy the bound $r^{2}\leq R$, giving further support for this conjecture. Moreover, we construct degree 2 maps with $r^{2}<R<1$ such that the target dimension $N$ is the embedding dimension. That means that there is no simple upper bound on $R$ if the degree $d$ is larger than 1.